Use the following probability distribution table to find the values for problems 2 3-30. 2 0.113 3 0.418 5 0.026 70.24!5 90.151 100.047
Answer #1

This is discrete probability distribution.

28)

P( X <= 5) = P( X = 2) + P( X = 3) + P( X = 5)

= 0.113 + 0.418 + 0.026

= 0.557

P( X <= 5) = 0.557

30)

P( 3 <= X < 7) = P( X = 3) + P( X = 5)

= 0.418 + 0.026

= 0.444

P( 3 <= X < 7) = 0.444